Mary is driving to San Francisco. Suppose that the remaining distance to drive (in miles) is a linear function of her driving time (in minutes). When graphed, the function gives a line with a slope of -0.9 .
Mary has 75 miles remaining after 50 minutes of driving. How many miles will be remaining after 64 minutes of driving?

Answer by htmentor(789)

Mary is driving to San Francisco. Suppose that the remaining distance to drive (in miles) is a linear function of her driving time (in minutes). When graphed, the function gives a line with a slope of -0.9 .
Mary has 75 miles remaining after 50 minutes of driving. How many miles will be remaining after 64 minutes of driving?
===============================================
The function will have the form d = m*t + d0
where d0 = the distance remaining at t=0 and m is the slope
Given: m = -0.9, d=75,t=50
Use this information to solve for d0:
75 = -0.9*50 + d0
d0 = 75 + 45 = 120
So the equation is d = -0.9t + 120
Solve for d when t=64:
d = -0.9*64 + 120
d = 62.4 minutes